### v2

@@ Abstract
Airdrops are used by blockchain applications and protocols to attract an initial user base, and to grow the user base over time. In the case of many airdrops, tokens are distributed to select users as a “reward” for interacting with the underlying protocol, with a long-term goal of creating a loyal community that will generate genuine economic activity well after the airdrop. Although airdrops are widely used by the blockchain industry, a proper understanding of the factors contributing to an airdrop’s success is generally lacking.
In this work, we outline the design space for airdrops, and specify a reasonable list of outcomes that an airdrop should ideally result in. We then analyze on-chain data from several larger-scale airdrops to empirically evaluate the success of previous airdrops, with respect to our desiderata. In our analysis, we demonstrate that airdrop farmers frequently dispose of the lion’s share of airdrops proceeds via exchanges. Our analysis is followed by an overview of common pitfalls that common airdrop designs lend themselves to, which are then used to suggest concrete guidelines for better airdrops.
keywords:
Airdrop design, Airdrop farmer, Web3, Layer 2
@@ 1 Introduction
Although the fundamental concept of airdrops is relatively straightforward, the design space for such reward schemes is expansive and may vary depending on the characteristics of the platform implementing the airdrop. For instance, certain airdrop mechanisms focus on power users, granting them a substantial amount of rewards with the expectation that these users will stimulate valuable economic activity, subsequently attracting additional users. However, a potential drawback to this approach arises when the distributed tokens empower users to propose changes to a protocol through decentralized governance. It allows users to suggest or vote on changes within a platform, typically employing a one-token-one-vote strategy, where a single user may possess more than one voting token. This approach introduces the risk of voting power concentration, where a minority of users hold the majority of the decision-making voting power [56, 22, 68].
@@
To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops. Armed with a basic desiderata, we then go over previous airdrops, see how they measure up, and uncover some interesting insights in the process. Specifically, we analyze the data pertaining to five popular airdrops (ENS [19], dYdX [39], 1inch [10], Arbitrum [41], and Uniswap [1]) and a fake airdrop conducted by a sybil farmer (Gemstone [35, 50]). We conclude that most of the proceeds (as high as 95%) have been sold via exchanges shortly after the airdrop, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers, highly-skilled users who employ sophisticated tactics to increase the share of airdropped tokens that they receive. Furthermore, we describe the common challenges faced by airdrops in the past. Given that the blockchain ecosystem is still quite young, it is expected that the theory and practice of blockchain airdrops are still in its infancy and thus that previous airdrops may not have been entirely successful in the long run. Finally, we use our analysis to suggest potential ways to create better airdrop mechanisms that are more fair to honest users.
1.1 Research Questions
In this section, we address three critical research questions central to our study.
[…]
We evaluate the effectiveness of blockchain protocols in converting short-term user incentives, such as airdrops, into long-term user engagement. The primary goal of conducting an airdrop is to attract users in the short-term and encourage them to become long-term users. Our investigation aims to measure how well these protocols achieve this objective.
Contrary to the assumption that airdrops are the primary driver of long-term user base growth, our research reveals otherwise. Through an exploration of past airdrops conducted, we demonstrate that airdrops may not be the sole determinant of sustained user retention. For example, findings from a recent report by Messari [8] indicate that users join a platform when short-term incentives are available, such as airdrops. However, these protocols struggle to retain users in the long-term, as users often shift their attention to protocols offering better incentives as soon as the initial rewards expire. This behavior implies that protocols successfully onboard users during incentive periods, like airdrops, but witness a drop in user engagement once the incentives come to an end.
Additionally, protocols not conducting airdrops also manage to onboard users due to an expectation of potential future airdrops. Our findings align with this observation, as our analysis indicates that, even after Arbitrum conducted an airdrop, other protocols attracted user attention and increased their total value locked (TVL). This suggests that airdrops alone may not be sufficient to ensure sustained user retention of the platform performing the airdrop, and other factors play crucial roles in long-term platform success.
[…]
Since protocols face challenges in achieving sustained user retention through airdrops alone, an alternative strategy involves incentivizing developers and innovators to build new applications, such as decentralized exchanges (DEXs)and non-fungible token (NFT). This approach aims to foster a dynamic ecosystem that benefits all users on the platform. To enhance user experience and maintain user loyalty, a focus on developing more, potentially exclusive, decentralized applications (dApps) is crucial. Another interesting approach is to provide additional incentives, such as discounts on fees for future platform interactions. For example, in the context of a rollup, users who receive airdropped tokens could redeem them for discounted fees, encouraging long-term engagement. Also, implementing multi-round airdrops allows protocols to learn from previous rounds, enabling iterative improvements to address previously identified issues [2]. Community engagement also plays a vital role in understanding the needs and preferences of users.
In line with these considerations, Blast has introduced an innovative airdrop strategy centered around user contributions to their protocol [55]. This strategy involves awarding points to users as a quantitative measure of their contribution. Points are granted for actions such as successfully bringing in new participants to the blockchain protocol, similar to a referral system, or for users who bridge (send) their tokens from one chain to Blast.
Furthermore, user engagement in blockchain protocols often hinges on short-term incentives. Users tend to join the platform when these incentives, such as airdrops, are available. However, a notable trend is the propensity of users to migrate opportunistically from one blockchain to another, following the trail of incentives. Consequently, a platform experiences an influx of users during incentive periods, only to witness a decline when these incentives cease.
[…]
A crucial aspect to consider involves airdrop farmers. These are sophisticated users who employ intricate tactics to maximize their chances of receiving more tokens per airdrop. They often create numerous wallet addresses controlled by a single entity, a practice known as Sybil [37, 64]. Detecting Sybils can be challenging [21, 51, 36, 73, 59], especially when these accounts execute automatic and randomized actions (e.g., token transfers, bridging, providing liquidity, exchanging tokens, NFTs, etc.) on the blockchain without linking their wallets. Furthermore, the accessibility of airdrop farming software facilitates the automated execution of such actions [70, 29, 12, 15]. Given its intricacy, current methods for identifying Sybils predominantly rely on graph network analysis [21, 51] and machine learning techniques [36, 73, 59].
Another common challenge involves airdropping governance tokens. These tokens empower holders to propose or vote on protocol changes. Distributing these tokens can be risky when users do not prioritize the protocol’s best interests. Additionally, it poses the danger of granting excessive power to certain users, as these protocols typically use a one-token-one-vote mechanism [56, 22, 68]. While acknowledging the challenges highlighted earlier, we believe that these challenges can serve as valuable lessons for protocols venturing into the realm of airdrops, offering insights into the potential path to success.
Furthermore, it is essential to uphold a positive reputation within the community by avoiding any occurrences of insider trading [5, 42]. Insider trading refers to individuals within an organization using privileged information to gain a financial edge over other protocol users. This practice is a common violation of securities laws in traditional financial markets [6], and engaging in it can lead to negative reactions from the blockchain community [69, 61, 23].
Finally, the points discussed above raise questions about whether protocols are effectively achieving their goals post their airdrop campaigns.
1.2 Summary of our Contributions
@@ 2 Airdrop Desiderata
Airdrops are powerful tools for promoting protocols and acquiring users, attracting newcomers, and incentivizing existing users to engage with the blockchain protocol or application. They are widely utilized for these purposes, with numerous examples documented in the literature (see Table 1). Any protocol capable of creating tokens can utilize airdrops to distribute them to users. For example, blockchain rollup solutions like Arbitrum and Optimism, as well as DeFi applications such as Uniswap, 1inch, dYdX, and ENS have all employed airdrops.
Airdrops can manifest in various forms, with single-round and multi-round distributions being the most prevalent [2]. In a single-round airdrop, tokens are distributed to users all at once, whereas multi-round airdrops involve distributing tokens over multiple rounds each following different strategies. This approach allows leveraging insights from previous rounds to address challenges encountered, such as mitigating potential sybil (i.e., multiple accounts controlled by a single entity) by observing past user behavior patterns [2]. The choice between single-round and multi-round airdrops depends on the protocol’s objectives and community dynamics. Community engagement is a critical aspect of airdrop development; it should be carefully considered by protocol developers.
Moreover, the timing of an airdrop significantly impacts the number of eligible users receiving tokens. Established protocols opting for delayed airdrops may have a larger user base compared to newer projects executing airdrops sooner. This difference in user base size introduces an added layer of complexity when performing sybil detection, as there are more accounts to evaluate and potentially safeguard against. If not performed correctly, this could impact community happiness. In such instances, the protocol may inadvertently reward accounts linked to industrial farming, which could be viewed negatively by the community. However, addressing industrial farming poses significant challenges and complexities. To mitigate this issue, LayerZero Labs implemented a self-reporting mechamism for sybils. Under this system, sybils have the option to self-report and receive 15% of their designated token allocation. Failure to self-report may result in them not receiving tokens if they are identified by the company or its affiliates [46].

### v3

@@ Abstract
Airdrops are a common strategy used by blockchain protocols to attract and grow an initial user base. Tokens are typically distributed to select users as a “reward” for engaging with the protocol, aiming to foster long-term community loyalty and sustained economic activity. Despite their prevalence, there is limited understanding of what makes an airdrop successful. This paper outlines the design space for airdrops and proposes key outcomes for an effective strategy. We analyze on-chain data from six large-scale airdrops to assess their success and find that a substantial portion of tokens is often sold off by “airdrop farmers.” Based on this analysis, we highlight common pitfalls and offer guidelines for improving airdrop design.
@@ 1 Introduction
Although the fundamental concept of airdrops is relatively straightforward, the design space for such reward schemes is expansive and may vary depending on the characteristics of the platform implementing the airdrop. For instance, certain airdrop mechanisms focus on power users, granting them a substantial amount of rewards with the expectation that these users will stimulate valuable economic activity, subsequently attracting additional users. However, a potential drawback to this approach arises when the distributed tokens empower users to propose changes to a protocol through decentralized governance, typically employing a one-token-one-vote strategy, where a single user may possess more than one voting token. This introduces the risk of voting power concentration, where a minority of users hold the majority of the decision-making power [17, 48, 60].
@@
To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops. We go over previous airdrops, see how they measure up, and uncover some interesting insights in the process, comparing them to a basic desiderata. Specifically, we analyze the data pertaining to five popular airdrops (ENS [14], dYdX [33], 1inch [7], Arbitrum [35], and Uniswap [1]) and a fake airdrop conducted by a Sybil farmer (Gemstone [30, 42]). We conclude that most of the proceeds (as high as 95%) have been sold via exchanges shortly after the airdrop, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers, highly-skilled users who employ sophisticated tactics to increase the share of tokens that they receive. Furthermore, we describe the common challenges faced by airdrops in the past. Given that airdrops are a relatively recent phenomenon, it is expected that the theoretical and practical understanding of them are still in their infancy, and thus previous airdrops may not have been entirely successful in the long run. Finally, we use our analysis to suggest ways to create better airdrop mechanisms that are more fair to honest users.
@@ 2 Airdrop Desiderata
Airdrops are powerful tools for promoting protocols and acquiring users, attracting newcomers, and incentivizing existing users to engage with these protocols and their applications. They are widely utilized for these purposes, with numerous examples documented in the literature (see Table 1). Protocols can create tokens and use airdrops to distribute them to users. For example, blockchain rollup solutions like Arbitrum, Optimism, and ZKsync Era, as well as DeFi applications such as Uniswap, 1inch, dYdX, and ENS have all employed airdrops.
Airdrops can manifest in various forms, with single-round and multi-round distributions being the most prevalent [2]. In a single-round airdrop, tokens are distributed to users all at once, whereas multi-round airdrops involve distributing tokens over multiple rounds each following different strategies. This approach allows leveraging insights from previous rounds to address challenges encountered, such as mitigating potential Sybil (i.e., multiple accounts controlled by a single entity) by observing past user behavior patterns [2]. The choice between single-round and multi-round airdrops depends on the protocol’s objectives and community dynamics.
Moreover, the timing of an airdrop significantly impacts the number of eligible users receiving tokens. Established protocols opting for delayed airdrops may have a larger user base compared to newer projects executing airdrops sooner. This difference in user base size introduces complexity when performing Sybil detection, as there are more accounts to evaluate and potentially safeguard against. If not performed correctly, this could impact community happiness. In such instances, the protocol may inadvertently reward accounts linked to industrial farming, which could be viewed negatively by the community. To mitigate this issue, LayerZero Labs implemented a self-reporting mechamism for sybils [39]. Under this system, sybils have the option to self-report and receive 15% of their designated token allocation.
@@
Airdrop Start date End date Blockchain Occurrence Project type
ENS [14] Oct. 2021 May 2022 Ethereum Single Name service
dYdX [33] Sept. 2021 Still active Ethereum Continuous DEX
1inch [7] Dec. 2020 Still active Ethereum Continuous DEX
@@
Arbitrum [35] Mar. 2023 Sept. 2023 Arbitrum One Single Single
Uniswap [1] Sept. 2020 Still active Ethereum Continuous DEX
Next, we decompose airdrop’s high-level goal to bootstrap a community of users into multiple sub-goals. These sub-goals are not necessarily orthogonal and there may be others that are worthwhile to pursue; we focus on them as they serve to illustrate interesting issues faced by common airdrop mechanisms.
Foster long-term involvement.  Bootstrapping an initial base of users is important, but not enough for sustaining a high degree of economic activity over time. Ideally, users should become regular, everyday users. This can be achieved by issuing rewards that can only be used within the blockchain protocol or application, similarly to frequent-flyer points. For example, in Layer-2 blockchains (L2s) [59], one may award fees discounts on future transactions. Other measures that may help are having multiple airdrops and rewarding users who perform “quests” that expose them to what the protocol has to offer. For example, Linea’s airdrop quests gave users an in-depth tour of their features and use-cases [37].
Target value-creating users.  An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of their platform. In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools [72] and Decentralized Exchanges (DEXs) [70], or across a variety of tokens. In rollups, one may consider so-called “creators” who deploy popular and useful contracts or users who bridge tokens to the rollup [74] as particularly valuable. Such users provide additional use cases for a platform, and thus serve to attract other users.
3 Exchange Aftermarket Analysis
The motivation for our quantitative analysis in this paper comes from the following observation: airdrop recipients quickly sell tokens and jump ship, which obviously defeats the point of airdrops in the first place. Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them. For example, shortly after the ParaSwap’s drop, 61% of their tokens where sold [16].
In both cases, the majority of recipients ceased using the relevant blockchain protocols within a few months. This pattern suggests that the airdrop was ineffective in keeping recipients engaged, or that many recipients were likely Sybil accounts. Moreover, a rapid sell-off could unsettle the market, particularly if seen as an indication of dwindling confidence in the protocol’s future prospects. Here, we examine data associated with six airdrops presented in Table 1, which were gathered from archive nodes of Ethereum, Arbitrum, and ZKsync Era (see §0.A for details on our data collection). To identify exchanges, we use a list of 620 exchange account addresses obtained from Dune and Etherscan [15].
Table 2: Airdrop distribution statistics for the six airdrops. Note that protocols often send a high share of airdropped tokens to addresses they control. For more details on the top earners, refer to Table 4 in §0.D.
Protocol Token Tokens % sold at # of recipient Tokens received per address
supply distributed exchanges addresses Mean Median Min Max
ENS $100\times 10^{6}$ $19.6\times 10^{6}$ 36.62% $102822$ $190.93$ $152.4$ $0.0001$ $1143.54$
dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.45% $42682$ $2325.85$ $392.56$ 00 $24.7\times 10^{6}$
1inch $1.5\times 10^{9}$ $83.2\times 10^{6}$ 54.05% $43402$ $1917.86$ $636.48$ 00 $9.75\times 10^{6}$
Gemstone $50.6\times 10^{15}$ $50.4\times 10^{15}$ 95.01% $20014$ $2.31\times 10^{12}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
Arbitrum $10\times 10^{9}$ $1.09\times 10^{9}$ 10.15% $583137$ $1874.02$ $1250$ $625$ $10250$
Uniswap $1\times 10^{9}$ $136.5\times 10^{6}$ 5.50% $220220$ $620.13$ $400$ $400$ $2.1\times 10^{6}$
3.1 Tokens Distributed Over Time
We provide quantitative insights into how airdropped tokens are disseminated across the user base in Table 2. Our analysis reveals a significant number of recipients for each airdrop, suggesting potential involvement of airdrop farmers (see column 5 in Table 2). Furthermore, the data indicate that tokens were frequently traded on exchanges. This is supported by the observation that the initial transfer made by the addresses post-airdrop was often a sale on an exchange (see column 4 in Table 2). Notably, the case of Gemstone is particularly striking, with 95% of its tokens being sold on exchanges. In this particular Gemstone case, a non-open sourced DEX created by the airdrop farmer [30].
(a) Daily token claims.
(b) Time to reach an exchange.
Figure 1: Comparison of token claiming and transfer patterns post-airdrop: (a) daily token claims; (b) time taken to reach an exchange.
Among all the airdrops analyzed in this study, the Arbitrum airdrop exhibited the fastest rate of token claims by users. Fig. 1(a) illustrates the distribution of accounts claiming their tokens each day. Arbitrum conducted a substantial airdrop, distributing $1162166000$ ARB tokens among $625143$ selected accounts. Notably, $583137$ (93.28%) of these accounts successfully claimed 94.03% of the allocated ARB tokens. Remarkably, 72.45% of the accounts completed their token claims on the first day, with an additional 14.41% doing so on the following day. This cumulative figure amounts to almost 87% of accounts claiming the Arbitrum airdrop within one day of its release, indicating high user engagement and prompt action among the majority of participants.
Table 3: Airdrop recipients’ interaction with exchanges. Gemstone and 1inch stands out as an exception with a median of 1 hop, while other protocols show a median of 2 hops for connecting airdrop recipients with exchanges.
Protocol Airdrop recipients # of hops to reach an exchanges
Total # sent to exchanges % sent to exchanges Mean Median Min. Max.
ENS $102822$ $86156$ 83.79% 1.65 2 1 50
dYdX $42682$ $40475$ 94.83% 1.77 2 1 17
1inch $43402$ $41913$ 96.57% 1.50 1 1 7
Gemstone $20014$ $20000$ 99.93% 1 1 1 1
Arbitrum $583137$ $555036$ 95.18% $2.34$ $2$ $1$ $31$
Uniswap $220220$ $213066$ 96.75% 2.21 2 1 17
Additionally, Table 3 also presents the shortest path from each recipient address of an airdrop to any exchange address in our dataset. Our findings reveal that the transfer of tokens to exchanges generally involves only a few steps, suggesting that airdrop recipients do not make significant efforts to conceal their activities. Gemstone is a notable outlier, with all tokens being sold within a single hop. Surprisingly, most accounts managed to reach an exchange with a relatively low number of intermediary hops (2 hops). This observation underscores the pivotal role that exchanges play within the cryptocurrency ecosystem.
The majority of accounts exchanged their tokens within approximately one million blocks post-airdrop. The number of blocks for Gemstone is significantly higher than others due to a delay introduced by its developer. Given the difference in block generation times between Ethereum (with a new block created every 15 seconds) and ZKsync, we normalized the blocks in terms of days. From Fig. 1(b), within a single day, 66.09% of the 1inch accounts interacted with an exchange. In comparison, ENS had an interaction rate of 55.15%, while dYdX stood at 64.26% followed by Arbitrum and Uniswap with 60.34% and 12.39%, respectively. This rapid exchange behavior does not align with one of the primary objectives of a protocol conducting an airdrop, which is to promote sustained user engagement, as discussed in §2. The prompt exchange of tokens suggests that users might be leaving a protocol soon after receiving their airdrops.
3.2 Token Distribution Graphs
To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network. Denoted as G(V, E), every node (V) represents an address, and an edge (E) is established when tokens are transferred from one address to another. Specifically, the ENS network consists of $184585$ nodes and $608462$ edges, the dYdX network comprises $112853$ nodes and $406027$ edges, the Gemstone network encompasses $20014$ nodes and $240113$ edges, and the 1inch network features $308329$ nodes and $1400913$ edges. Lastly, Arbitrum accounts for $2025898$ nodes and $27438608$ edges while Uniswap contains $1180830$ nodes and $3762613$ edges.
To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdrop the first few hours after the airdrop and plot their largest component in Fig. 7 in §0.B. We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer [15]. We show that for all protocols but Gemstone, the DEXthat received the majority of transfers (measured in terms of in-degree) from airdrop recipients is Uniswap, followed by SushiSwap.
In the case of Gemstone, all tokens were sent to the address 0x7$\cdots$49ad in just a single hop. On the other hand, in dYdX, the airdrop utilized a more diverse set of exchange addresses. Notably, as depicted in Table 3, a portion of the airdrop recipients chose to sell their tokens on an exchange. We see some common exchanges like Uniswap, Wintermute, and SushiSwap.
(a) Absolute number.
(b) Relative daily active addresses.
(a) Daily median fees.
(b) Daily average fees.
Figure 3: Daily transaction fees in USD. (a) Median; and (b) Average.
4 Measuring the Airdrop Lift
Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially [16]. Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic [16, 25]. In this section, we examine this through the lens of some relevant metrics such as daily transaction count, daily active addresses, median transaction cost, Total Value Locked (TVL), fees paid by users, and stablecoin market cap applied to the Arbitrum airdrop. We rely on data source from Growthepie [29]. For more details on this data refer to §0.A in the appendix.
4.1 Unique Active Addresses
Others outperformed Arbitrum without an airdrop.  While the number of unique addresses on Arbitrum increased post-airdrop (March 22nd 2023 [57]) and remained 50% higher than pre-airdrop levels, other protocols achieved similar growth without an airdrop. For example, per Fig. 2(b), Optimism saw a larger increase in addresses in May 2023, likely due to the Bedrock launch [54]. Similarly, ZKsync Era surpassed Arbitrum’s address count two months after the airdrop, as shown in Fig. 2(a).
Fees may explain the closing gap between Arbitrum and Optimism.  Arbitrum has consistently led Optimism in daily active addresses. However, Fig. 8 in §0.C shows this gap narrowing. Before the airdrop, Arbitrum had 2.6 times more active addresses than Optimism, but in the last 50 days, this has decreased to 1.83 times. Optimism’s lower median transaction fees since June may partly explain this trend (Fig. 3(a)), with the average fee gap between the two protocols being even smaller (Fig. 3(b)).
Unique address counts oscillates over time.  The number of unique addresses shows oscillatory behavior, as seen in Fig. 2(b), with rapid rises, peaks, and drops. Notably, Optimism and Arbitrum display inverted phases in relative address counts, possibly due to users switching between protocols as fees rise. However, Fig. 3(a) shows no such pattern in median transaction fees, with Arbitrum consistently having lower fees in mid-May 2023. However, unique address counts metric can be gameable. Per our analysis, Arbitrum’s airdrop did not lead to long-lasting user involvement, as the unique address metric can be gamed. Users can create multiple addresses (Sybil attacks) to exploit airdrop limits. This makes the metric unreliable for measuring genuine activity, as large fluctuations may result from such behavior. Furthermore, the accessibility of airdrop farming software facilitates the automated execution of such activities [10, 24, 53, 63]. Thus, alternative, more Sybil-resistant metrics should be considered to gauge authentic user involvement such as graph network analysis [16, 43] combined with machine learning techniques [31, 50, 66].
4.2 Transaction-related Metrics
Transaction-related metrics provide a useful proxy for measuring “real” economic activity, as users must pay a fee to send transactions [27], excluding cases with retroactive rebates or protocols operating at a loss [55].
The Arbitrum-Optimism gap narrows when considering transactions.  Notably, the transaction count gap between Arbitrum and Optimism had nearly closed by the end of July (see Fig. 4(a)). Additionally, Fig. 4(b) shows that Immutable X’s daily transaction count has almost halved since Arbitrum’s airdrop, whereas its unique address count remained relatively stable. This suggests that while address numbers held steady, user engagement with Immutable X has decreased.
(a) Absolute daily transactions.
(b) Relative daily transactions.
Arbitrum’s per-address transaction count declined post-airdrop.  To assess how Arbitrum’s airdrop impacted user engagement in other protocols, Fig. 5(a) shows the relative average daily transaction count per unique address. Since the airdrop, Arbitrum’s per-user transaction count has dropped to less than 75% of its pre-airdrop level. However, since transaction counts can be misleading without considering fees, high transaction counts alone may not indicate genuine engagement. In this regard, some protocols require users to perform multiple transactions to qualify for airdrops, leading to inflated activity when fees are low. This aligns with Goodhart’s law [46], which states that “when a measure becomes a target, it ceases to be a good measure”.
(a) Daily transactions per address.
(b) Average fees spent per address.
Figure 5: Daily transactions per address (a) and daily average fees per address (b), relative to the average before Arbitrum’s airdrop.
Since June, average transaction fees are similar for all protocols.  Good metrics should reflect a user’s commitment, and transaction fees can serve as a proxy for this since it measures how much a user is willing to expend for interacting with the protocol. Since June, the average fee per transaction and per unique address have been similar across protocols. Comparing average and median fees (Fig. 3) shows that the median fee may be more useful in understanding shifts in user behavior. Another useful metric is the relative average fees per address compared to the 50 days before Arbitrum’s airdrop, as shown in Fig. 5(b). It shows that Arbitrum’s user engagement was not significantly impacted by the airdrop and generally follows similar patterns to other protocols.
Arbitrum’s total daily fees spiked around the time of the airdrop.  Arbitrum’s airdrop did not provide a significant long-term advantage over its competitors in terms of income from transaction fees as shown in Fig. 9 in §0.C. While Arbitrum experienced a peak in fees on the day of the airdrop, this increase was short-lived. Indeed, per Fig. 6(a), in the 50 days before the airdrop, Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
(a) Ratio of daily transaction fees.
(b) TVL growth.
Figure 6: Comparison of the ratio of daily transaction fees between Arbitrum and Optimism (a); and the growth in Total Value Locked (TVL) relative to the average of the 50 days before Arbitrum’s airdrop (b).
4.3 Total Value Locked (TVL)
The TVLof a protocol is a metric that measures the combined value of all assets stored within that protocol.